# Fine-Tuned Language Models Generate Stable Inorganic Materials as Text

### v1

@@ 1 Introduction
Large language models (LLMs) are trained to compress large text datasets, but can also act as strong foundations for non-text data (Delétang et al., 2023). As compressors, LLMs extract common patterns and find simple programs that can produce them (Goldblum et al., 2023; Sutskever, 2023), regardless of the data’s origin. From text pretraining alone, LLMs can compress or extrapolate data as diverse as images (Delétang et al., 2023), tabular data (Goldblum et al., 2023), time series (Gruver et al., 2023a), or robotic trajectories (Mirchandani et al., 2023). Alongside generality, LLM pre-training also gives rise to sample efficiency, as in-context learning and fine-tuning require far fewer training examples to identify salient patterns than training a model from scratch (Brown et al., 2020).
@@ 2 Related Work
Xie et al. (2021) introduced crystal diffusion variational autoencoder (CDVAE) to directly deal with both of these challenges. CDVAE uses several individual generative models for discrete and continuous components that share a continuous (VAE) latent space. The chemical composition is reconstructed from this latent space using a language modeling head, while atom positions are generated with a denoising diffusion model (Ho et al., 2020). Since CDVAE, several works have extended diffusion processes to capture all parameters of the crystal, not just the atomic coordinates. Both Jiao et al. (2023) and Zeni et al. (2023) accomplish this by creating diffusions for the lattice parameters and atom identities, while Yang et al. (2023) design a new continuous representation that unifies atom identities and positions in a single high-dimensional tensor. In most cases, these diffusion models were designed with a careful eye towards symmetries and are built on top of graph neural networks with strict invariance/equivariance properties (Xie et al., 2021; Jiao et al., 2023; Zeni et al., 2023). The approach of Yang et al. (2023) is more similar to ours, as they apply a general-purpose architecture (3D U-net) and modeling approach (Gaussian diffusion) to a new representation, without guaranteeing symmetries. Discrete atom identities and variable length (number of atoms), however, require special considerations in diffusion models, unlike standard language models, which were originally designed for modeling discrete sequences.
@@ 3 Background
Stability of hypothetical materials ($E_{hull}$)    The composition of a crystal also impacts its energy, as different elements have different geometries and charge properties. Certain stoichiometries, or ratios of elements, are naturally favored, and a composition of elements $A$ and $B$ with constituent parts $A_{x}B_{y}$ can dissociate into the composition $A_{c}B_{d}$ if it is energetically favorable. Because of the effect of composition, the energy of a crystal is typically a two dimensional concept captured by the energy hull, which is the minimum observed configuration energy for a given composition. For a crystal to be low-energy and stable, and therefore give rise to a practically useful material, it must have a small energy above $E_{\text{hull}}$), the distance from the energy hull for the crystals elemental composition (Figure 2). Crystals with $E_{\text{hull}}<0$ are considered stable and by definition have lower energy than the known minimum (which has $E_{\text{hull}}=0$). Crystals with $E_{\text{hull}}<0.1$ eV/atom are often metastable and likely to be practical useful (Sun et al., 2016).
@@ 4 Method
Crystals structures are symmetric under translational. All atomic coordinates can be shifted modulo the lattice boundaries without changing the resulting material structure. Similarly, the ordering of atoms within the lattice is irrelevant to the underlying material (permutation invariance). Prior work on diffusion generative models guarantee these symmetries as invariance or equivariance constraints on the model architecture (Xie et al., 2021; Jiao et al., 2023). To encourage translation invariance in our language models, we apply random uniform translations to the fractional coordinates. We chose not to augment the ordering of atoms because these variables often contained valuable information, for example grouping set of elements together for placement in the lattice (discussion in Appendix A.1).
@@ 5 Experiments
For consistency with prior work (Xie et al., 2021; Flam-Shepherd et al., 2023) we used MP-20 (Jain et al., 2013), a dataset of 45231 materials, when training for unconditional generation. All structures in MP-20 are stable, and therefore an effective generative model trained on MP-20 should tend to propose new crystals that are at least metastable. For text-conditioned generation, we train with all forms of prompting (Section 4) on a collection of  120,000 crystals from Materials Project (Appendix A.3). The collection includes basic property information, such as the space group number, band gap, $E_{\text{hull}}$ and the chemical formula. All of our experiments were conducted with LLaMA-2 models (7B 13B, and 70B) (Touvron et al., 2023a; b) through the Transformers library (Wolf et al., 2020) and PyTorch (Paszke et al., 2019). In order to train on small number of GPUs we use 4-bit quantization (Dettmers et al., 2022) and Low-Rank Adapters (LoRA) (Hu et al., 2021). We provide the full hyperparameters and training details in Appendix A.4.
@@
For basic evaluation of the LLM samples, we use the validity and diversity metrics introduced by Xie et al. (2021). Structural validity is determined by non-overlapping atomic radii (overlapping taken to be both atoms within half a radius of each other), while compositional validity captures the net charge of the structure (only structures with net neutral total charge are valid). Diversity is computed as pairwise distance between samples under featurizations of the structure and composition from Matminer (Ward et al., 2018; Xie et al., 2021).
@@
${}^{\dagger}$ Fraction of structures that are first predicted by M3GNet to have $E_{\text{hull}}^{\text{M3GNet}}<0.1$ eV/atom, and then verified with DFT to have $E_{\text{hull}}^{\text{DFT}}<0.0$ eV/atom.
@@
In many practical settings, sampling and filtering materials from scratch is unnecessary. Good starting materials are often known, and manufacturing processes are easier to adapt to related compositions than develop completely from scratch by making small edits to their composition–often referred to as template methods (Kirklin et al., 2015; Saal et al., 2013). To emulate a typical template method, we construct a lookup table that maps each element to elements that have a similar atom radius when in the same oxidation state (code in Appendix C). We choose an element uniformly at random and swap it with a random element chosen from the table. The resulting structure is then relaxed using M3GNet. To improve this strategy using our fine-tuned LLM, we used the infilling prompt (Section 4) to obtain a distribution over elements (modulated with temperature $\tau$) which we use instead of a uniform distribution over swaps. To evaluate our mutation procedure, we sample 3000 structures randomly from the test set and generate perform one mutation-relaxation step for each, using both uniform and language model-guided sampling. In Figure, 7 we show the percentage of stable compounds and diversity in the stable compounds for the uniform baseline and LLaMA-2 70B with different temperature values. LLaMA-2 70B proposes elements that lead to stable structures at a higher rate than the baseline template method without sacrificing diversity.
@@ 6 Discussion
Future directions    There is substantial room for improvement in conditional generation, which could be used to directly generate materials with desired properties. While we did not pursue alternative sampling strategies in depth, approaches like classifier-free guidance (Sanchez et al., 2023) or variants of PPLM (Dathathri et al., 2019) might be useful in combination with fine-tuned LLMs to improve conditional generation. These methods could also be combined with primitives from Bayesian optimization for sample-efficient and uncertainty-aware design (Stanton et al., 2022; Gruver et al., 2023b).

### v2

@@ 1 Introduction
Large language models (LLMs) are trained to compress large text datasets, but can also act as strong foundations for non-text data (Delétang et al., 2023). As compressors, LLMs extract common patterns and find simple programs that can produce them (Goldblum et al., 2023, Sutskever, 2023), regardless of the data’s origin. From text pretraining alone, LLMs can compress or extrapolate data as diverse as images (Delétang et al., 2023), tabular data (Goldblum et al., 2023), time series (Gruver et al., 2023a), or robotic trajectories (Mirchandani et al., 2023). Alongside generality, LLM pre-training also gives rise to sample efficiency, as in-context learning and fine-tuning require far fewer training examples to identify salient patterns than training a model from scratch (Brown et al., 2020).
@@ 2 Related Work
Xie et al. (2021) introduced crystal diffusion variational autoencoder (CDVAE) to directly deal with both of these challenges. CDVAE uses several individual generative models for discrete and continuous components that share a continuous (VAE) latent space. The chemical composition is reconstructed from this latent space using a language modeling head, while atom positions are generated with a denoising diffusion model (Ho et al., 2020). Since CDVAE, several works have extended diffusion processes to capture all parameters of the crystal, not just the atomic coordinates. Both Jiao et al. (2023) and Zeni et al. (2023) accomplish this by creating diffusions for the lattice parameters and atom identities, while Yang et al. (2023) design a new continuous representation that unifies atom identities and positions in a single high-dimensional tensor. In most cases, these diffusion models were designed with a careful eye towards symmetries and are built on top of graph neural networks with strict invariance/equivariance properties (Xie et al., 2021, Jiao et al., 2023, Zeni et al., 2023). The approach of Yang et al. (2023) is more similar to ours, as they apply a general-purpose architecture (3D U-net) and modeling approach (Gaussian diffusion) to a new representation, without guaranteeing symmetries. Discrete atom identities and variable length (number of atoms), however, require special considerations in diffusion models, unlike standard language models, which were originally designed for modeling discrete sequences.
@@ 3 Background
Stability of hypothetical materials ($E_{hull}$)    The composition of a crystal also impacts its energy, as different elements have different geometries and charge properties. Certain stoichiometries, or ratios of elements, are naturally favored, and a composition of elements $A$ and $B$ with constituent parts $A_{x}B_{y}$ can dissociate into the composition $A_{c}B_{d}$ if it is energetically favorable. Because of the effect of composition, the energy of a crystal is typically a two dimensional concept captured by the energy hull, which is the minimum observed configuration energy for a given composition. For a crystal to be low-energy and stable, and therefore give rise to a practically useful material, it must have a small energy above hull ($E_{\text{hull}}$), the distance from the energy hull for the crystals elemental composition (Figure 2). Crystals with $E_{\text{hull}}<0$ are considered stable and by definition have lower energy than the known minimum (which has $E_{\text{hull}}=0$). Crystals with $E_{\text{hull}}<0.1$ eV/atom are often metastable and likely to be practical useful (Sun et al., 2016).
@@ 4 Method
Crystals structures are symmetric under translational. All atomic coordinates can be shifted modulo the lattice boundaries without changing the resulting material structure. Similarly, the ordering of atoms within the lattice is irrelevant to the underlying material (permutation invariance). Prior work on diffusion generative models guarantee these symmetries as invariance or equivariance constraints on the model architecture (Xie et al., 2021, Jiao et al., 2023). To encourage translation invariance in our language models, we apply random uniform translations to the fractional coordinates. We chose not to augment the ordering of atoms because these variables often contained valuable information, for example grouping set of elements together for placement in the lattice (discussion in Appendix A.1).
@@ 5 Experiments
For consistency with prior work (Xie et al., 2021, Flam-Shepherd et al., 2023) we used MP-20 (Jain et al., 2013), a dataset of 45231 materials, when training for unconditional generation. All structures in MP-20 are stable, and therefore an effective generative model trained on MP-20 should tend to propose new crystals that are at least metastable. For text-conditioned generation, we train with all forms of prompting (Section 4) on a collection of  120,000 crystals from Materials Project (Appendix A.3). The collection includes basic property information, such as the space group number, band gap, $E_{\text{hull}}$ and the chemical formula. All of our experiments were conducted with LLaMA-2 models (7B 13B, and 70B) (Touvron et al., 2023a; b) through the Transformers library (Wolf et al., 2020) and PyTorch (Paszke et al., 2019). In order to train on small number of GPUs we use 4-bit quantization (Dettmers et al., 2022) and Low-Rank Adapters (LoRA) (Hu et al., 2021). We provide the full hyperparameters and training details in Appendix A.4.
@@
For basic evaluation of the LLM samples, we use the validity and diversity metrics introduced by Xie et al. (2021). Structural validity is determined by non-overlapping atomic radii (overlapping taken to be both atoms within half a radius of each other), while compositional validity captures the net charge of the structure (only structures with net neutral total charge are valid). Diversity is computed as pairwise distance between samples under featurizations of the structure and composition from Matminer (Ward et al., 2018, Xie et al., 2021).
@@
Fraction of structures that are first predicted by M3GNet to have $E_{\text{hull}}^{\text{M3GNet}}<0.1$ eV/atom, and then verified with DFT to have $E_{\text{hull}}^{\text{DFT}}<0.0$ eV/atom.
@@
In many practical settings, sampling and filtering materials from scratch is unnecessary. Good starting materials are often known, and manufacturing processes are easier to adapt to related compositions than develop completely from scratch by making small edits to their composition–often referred to as template methods (Kirklin et al., 2015, Saal et al., 2013). To emulate a typical template method, we construct a lookup table that maps each element to elements that have a similar atom radius when in the same oxidation state (code in Appendix C). We choose an element uniformly at random and swap it with a random element chosen from the table. The resulting structure is then relaxed using M3GNet. To improve this strategy using our fine-tuned LLM, we used the infilling prompt (Section 4) to obtain a distribution over elements (modulated with temperature $\tau$) which we use instead of a uniform distribution over swaps. To evaluate our mutation procedure, we sample 3000 structures randomly from the test set and generate perform one mutation-relaxation step for each, using both uniform and language model-guided sampling. In Figure, 7 we show the percentage of stable compounds and diversity in the stable compounds for the uniform baseline and LLaMA-2 70B with different temperature values. LLaMA-2 70B proposes elements that lead to stable structures at a higher rate than the baseline template method without sacrificing diversity.
@@ 6 Discussion
Future directions    There is substantial room for improvement in conditional generation, which could be used to directly generate materials with desired properties. While we did not pursue alternative sampling strategies in depth, approaches like classifier-free guidance (Sanchez et al., 2023) or variants of PPLM (Dathathri et al., 2019) might be useful in combination with fine-tuned LLMs to improve conditional generation. These methods could also be combined with primitives from Bayesian optimization for sample-efficient and uncertainty-aware design (Stanton et al., 2022, Gruver et al., 2023b).
Acknowledgements. We thank NSF CAREER IIS-2145492, NSF HDR-2118310, and NSF CDS&E-MSS 2134216 for support.